lcm:12 1/3 - 8 4/5 is the ans. 3 3/15
17 4/12 - 10 3/4 is the ans. 6 5/12

Answer by stanbon(48539)

lcm:12 1/3 - 8 4/5 is the ans. 3 3/15
Rewrite: 37/3 - 44/5
Then lcm = 15
Rewrite: (5*37)/lcm - (3*44)/lcm
= (185 + 132)/lcm
= 317/15
---------------------------
17 4/12 - 10 3/4 is the ans. 6 5/12
Rewrite: (12*17 + 4)/12 - (43/4)
lcm = 12
Rewrite: 208/12 - (3*43)/12
= (208 - 129)/12
= 79/12
===================================
Cheers,
stan H.